Suppose that the price of President’s Choice macaroni and cheese decreased from $10 to $8 per case, and at the same time, the quantity of Kraft macaroni and cheese sold dropped from 192 to 128 cases.



a) What is the cross-elasticity of demand between the two products?
1
Expert's answer
2018-05-24T09:36:08-0400
P1 = $10, P2 = $8 per case, Q1 = 192, Q2 = 128 cases.
a) The cross-elasticity of demand between the two products is:
Ed = (Q2 - Q1)/(P2 - P1)*(P2 + P1)/(Q2 + Q1) = (128 - 192)/(8 - 10)*(8 + 10)/(128 + 192) = 64/2*18/320 = 1.8, so these products are substitutes.